OpenStudy (anonymous):

Find tan2θ if tanθ = -1/2 -4/3 -1 -4/5 -3/4

OpenStudy (anonymous):

@Tommynaiter

OpenStudy (anonymous):

i have no idea how to do any of it

Nnesha (nnesha):

do you know sum and difference identities ?

Nnesha (nnesha):

I mean double angle formulas ?

OpenStudy (anonymous):

nope i wasnt taught any of this and im over half way done with the course

Nnesha (nnesha):

 according to double angle formulas \[\large\rm \tan(2 \theta)=\frac{2 \tan \theta}{1-\tan^2 \theta}\]

Nnesha (nnesha):

here is good link with all identities http://www.analyzemath.com/trigonometry/trigonometric_formulas.html you need to familiar with these in order to do trig :=))

Nnesha (nnesha):

now just use the given information tan theta=-1/2 plug that in

OpenStudy (anonymous):

im so confused

Nnesha (nnesha):

find \[\large\rm \tan2θ \] use the double angle formula `tan2θ ` can be written as \[\frac{2 \tanθ }{1-\tan^2θ }\] and since tan theta =-1/2 you can replace all tan with -1/2